(3,5,7,9).(3,5,7,9).


[[ Set B ]]

  1. Find ALL possible samples of size 2 which can be drawn with replacement from the population.
  2. Solve for the mean μ
  3. x
  4. μx of the sampling distribution of means.
  5. Solve for the variance σ
  6. x
  7. 2
  8. σx2 of the sampling distribution of means.
  9. Solve for the standard deviation σ
  10. x
  11. σx of the sampling distribution of means.

Round all answers to the nearest hundredths.


1
Expert's answer
2022-05-05T10:10:31-0400

We have population values 3,5,7,9, population size N=4 and sample size n=2.

Mean of population "(\mu)" = 

"\dfrac{3+5+7+9}{4}=6"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{9+1+1+9}{4}=5"

"\sigma=\sqrt{\sigma^2}=\sqrt{5}\approx2.24"

The number of possible samples which can be drawn without replacement is "2^4=16."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 3,3 & 3 \\
 \hdashline
 2 & 3,5 & 4 \\
 \hdashline
 3 & 3,7 & 5 \\
 \hdashline
 4 & 3,9 & 6 \\
 \hdashline
 5 & 5,3 & 4 \\
 \hdashline
 6 & 5,5 & 5 \\
 \hdashline
 7 & 5,7 & 6 \\
 \hdashline
 8 & 5,9 & 7 \\
 \hdashline
 9 & 7,3 & 5 \\
 \hdashline
 10 & 7,5 & 6 \\
 \hdashline
 11 & 7,7 & 7 \\
 \hdashline
 12 & 7,9 & 8 \\
 \hdashline
 13 & 9,3 & 6 \\
 \hdashline
 14 & 9,5 & 7 \\
 \hdashline
 15 & 9,7 & 8 \\
 \hdashline
 16 & 9,9 & 9 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 3 & 1/16 & 3/16 & 9/16\\
 \hdashline
 4 & 2/16 & 8/16 & 32/16\\
 \hdashline
 5 & 3/16 & 15/16 & 75/16\\
 \hdashline
 6 & 4/16 & 24/16 & 144/16\\
 \hdashline
 7 & 3/16 & 21/16 & 147/16\\
 \hdashline
 8 & 2/16 & 16/16 & 128/16\\
 \hdashline
 9 & 1/16 & 9/16 & 81/16\\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=6=\mu"


The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{616}{16}-(6)^2=\dfrac{5}{2}=2.5= \dfrac{\sigma^2}{n}"

"\sigma_{\bar{X}}=\sqrt{\dfrac{5}{2}}\approx1.58"